One of two complementary angles measures 30° more than three times the other. What is the measure of BOTH angles?
Answer by venugopalramana(3286)

let the 2 angles be x and y .
they are complementary hence x+y = 90 .......(1)
it is given that one say x measures 30 degrees more than 3 times the other that is y.
three times y = 3y...
30 degree more makes it 3y+30..
this is equal to x.
hence we have x=3y+30.....(2)..
substituting this in eqn.(1), we get
3y+30+y = 90
4y = 90-30 = 60
y=60/4 =15
but x+y = 90 so x=90-y=90-15 = 75
hence x=75 and y = 15.